A metal rod of length 40.0 cm at 20°C is heated through a temperature of 45°C if the new length of the rod

is 40.05 cm calculate its linear expansivity


1
Expert's answer
2022-02-20T15:44:46-0500

"\alpha=\frac{\Delta l}{l\Delta t}=5\cdot 10^{-5}~°C^{-1}."